glizxzy:

3.7x + 1.2 = 5.4x - 6.3

glizxzy:

@Salmon @sylvester @Aqual

Sweetie05:

I was tying to solve it but i dont think it is right

glizxzy:

do it i know the answer

glizxzy:

close

glizxzy:

@foxey3 @Salmon @Sacredcheats

foxey3:

yes

glizxzy:

solve it

foxey3:

k

sylvester:

im slow dont ask me 🥱

glizxzy:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sylvester im slow dont ask me 🥱 \(\color{#0cbb34}{\text{End of Quote}}\) cap

sylvester:

nope

glizxzy:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sylvester nope \(\color{#0cbb34}{\text{End of Quote}}\) yep

Aqual:

9.1x = 7.5

glizxzy:

nope

glizxzy:

?

Aqual:

ima figure it out

glizxzy:

ok

glizxzy:

@Pixel

Aqual:

x = -0.82417582417

glizxzy:

nerp

Aqual:

hmm

glizxzy:

answer?

Aqual:

ooooh, i know what i did wrong

Aqual:

-2.3x = -7.5

glizxzy:

nope

mxddi3:

i think i did this wrong lol. i got x=4.41176470588 im lost lmao

Aqual:

3.26086956522

glizxzy:

yep wrong

glizxzy:

nope

Aqual:

5

glizxzy:

yep

Aqual:

really?

Salmon:

x=4.41

glizxzy:

3.7x + 1.2 = 5.4x - 6.3 If the decimals are problematic, eliminate them by multiplying both sides of the equation by 10 10(3.7x + 1.2) = 10(5.4x - 6.3) 37x + 12 = 54x - 63 subtract 37x from both sides 37x - 37x + 12 = 54x - 37x - 63 12 = 17x - 63 Add 63 to both sides 12 + 63 = 17x - 63 + 63 75 = 17x divide both sides by 17 75/17 = 17x/17 5 = x

sylvester:

damn to many numbers 😷

glizxzy:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sylvester damn to many numbers 😷 \(\color{#0cbb34}{\text{End of Quote}}\) lol